# Causal Effect Estimation under Networked Interference without Networked Unconfoundedness Assumption

### v1

@@ 6.2 Experimental Analyses
Effectiveness of Our Method As shown in Table 1 and Table 2, we have conducted experiments by running our proposed methods and several baselines. Overall, our methods outperform all methods consistently with smaller estimation errors in all metrics, indicating the effectiveness of our methods. Specifically, compared with the baselines, our methods perform better in terms of both average and heterogeneous treatment effect estimation, with smaller errors and standard deviation. This is reasonable since existing methods do not consider the latent confounders that hinder their identifications. Both Ours and Ours_w/o_IPM are based on identifiable representation learning techniques and thereby achieve superior performances with recovered latent confounders. Compared ours with its variant without the IPM term, we found that the IPM term can slightly improve the performance. This is due to the fact that the IPM term effectively mitigates the confounding bias with balanced representations.
$\hat{u}^{n}$ $\hat{u}^{c}$ $\hat{u}^{i}$
${u}^{i}$ 0.2489 0.2775 0.9505
${u}^{c}$ 0.2179 0.8930 0.5070
${u}^{n}$ 0.9435 0.4125 0.1412
@@
Table 3: MCC results of recovered latent confounders.
Correctness of Representation Learning To validate the correctness of our representation learning method, We conduct experiments in the simulated dataset and visualize the recovered latent confounders $\hat{U}^{i},\hat{U}^{c},\hat{U}^{n}$ with ground-true $U^{i},U^{c},U^{n}$ in Figure 4. And we calculate the MCC results in Table 3. The result shows that the recovered latent confounders are highly correlated with the ground-true latent confounders, with very high MCC values. This indicates that our method can correctly recover the latent confounders, which validates the correctness of Theorem 4.2.
7 Conclusion
In this paper, we address the problem of networked causal effect identification and estimation in the presence of latent confounders. We leverage the networked information to achieve the identifiability of latent confounders. With identified latent confounders, we theoretically establish the identification result of networked effects. We further devise an effective estimator built on the theoretical findings. Extensive experiments validate the correctness of our theories and the effectiveness of our proposed estmator.
Impact Statement
This paper presents work whose goal is to advance the field of causal inference under networked interference. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Proof of Theorem 4.2
The proof can be directly obtained from the proof of Theorem 1 in Lu et al. (2022), with the following modifications:
  • we consider $X_{\mathcal{N}}$ as the prior conditional set;
  • we specify that the latent confounder $U$ contains three parts, i.e., $U=[U^{i},U^{c},U^{n}]$.
Appendix B Proof of Theorem 4.3
Proof.
Under Theorem 4.2, we can recover the whole distribution $p(x,u^{i},u^{c},u^{n},t,z,y)$, then the conditional average dose-response function is identified by
$\mu(t,z,x,x_{\mathcal{N}})$ $=\mathbb{E}[Y(t,z)|X=x,X_{\mathcal{N}}=x_{\mathcal{N}}]$ (12)
$=\mathbb{E}[\mathbb{E}[Y(t,z)|X=x,X_{\mathcal{N}}=x_{\mathcal{N}},U^{i}=u^{i},U^{c}=u^{c},U^{c}_{\mathcal{N}}=u^{c}_{\mathcal{N}},U^{n}_{\mathcal{N}}=u^{n}_{\mathcal{N}}]]$
$=\mathbb{E}[\mathbb{E}[Y(t,z)|X=x,X_{\mathcal{N}}=x_{\mathcal{N}},U^{i}=u^{i},U^{c}=u^{c},U^{c}_{\mathcal{N}}=u^{c}_{\mathcal{N}},U^{n}_{\mathcal{N}}=u^{n}_{\mathcal{N}},T=t,Z=z]]$
$=\mathbb{E}[\mathbb{E}[Y|X=x,X_{\mathcal{N}}=x_{\mathcal{N}},U^{i}=u^{i},U^{c}=u^{c},U^{c}_{\mathcal{N}}=u^{c}_{\mathcal{N}},U^{n}_{\mathcal{N}}=u^{n}_{\mathcal{N}},T=t,Z=z]]$
where the third equality is based on Assumption 3.11 and the forth equality is based on Assumption 3.7. and then the networked effect $\psi(t,z):=\mathbb{E}[Y(t,z)]$ is immediately identified. ∎
Appendix C Additional Experimental Details
C.1 Baseline Methods
The compared baselines in this paper are
  • TARNET+z: Original Tarnet (Johansson et al., 2021) uses two-heads neural networks, serving as T-Learner-like estimator, to estimate causal effects under no interference assumption. We modify TARNET by additionally inputting the exposure $z_{i}$.
  • CFR+z: Original CFR (Johansson et al., 2021) uses two-heads neural networks with an MMD term to achieve counterfactual regression under no interference assumption. We modify CFR by additionally inputting the exposure $z_{i}$.
  • ND+z: Original ND (Guo et al., 2020) propose network deconfounder framework by using network information under no interference assumption. We modify ND by additionally inputting the exposure $z_{i}$.
  • GEst (Ma & Tresp, 2021): GEst, based on CFR, uses GCN to aggregate the features of neighbors and input the exposure $z_{i}$ to estimate causal effects under networked interference.
  • NetEst (Jiang & Sun, 2022): NetEst learns balanced representation via adversarial learning for networked causal effect estimation.
  • TNet (Chen et al., 2024): TNet utilizes targeted learning techniques into its neural networks model to estimate causal effects under networked interference in a double robust manner.
C.2 Detailed Data Generation of Synthetic Dataset
We first generate the network graph $E$ with expected degrees $5$.
According to $E$, we directly generate all samples’ $u^{i},u^{c},u^{n}$ from a multivariate Gaussian distribution. We set the sample size as $1000$, then we first sample $\mathbf{u}=[[u^{i}]^{T},[u^{c}]^{T},[u^{n}]^{T}]^{T}$ as a $3000$-dimensional vector. To mimic the correlation among $u^{i},u^{c},u^{n}$ and between units, we generate the $3000\times 3000$ covariance matrix $V$ with all elements equal to $0.1$ besides diagonal equal to $1$. This results in that, for the same units, $u^{i}_{i},u^{c}_{i},u^{n}_{i}$ are correlated with covariance $0.1$, and for the different units, their latent confounders are also correlated with covariance $0.1$, e.g. $u_{i}^{i}$ and $u_{i}^{c}$. Then we generate $\mathbf{u}\sim\mathcal{N}(\mathbf{1},V)$ and reshape $\mathbf{u}$ as $[u^{i},u^{c},u^{n}]$, i.e., reshape to obtain the $1000\times 3$ design matrix of $u$.
Then, given $[u^{i},u^{c},u^{n}]$, we generate treatment $t$ as
$t_{i}\sim\textbf{Bern}(1/(1+\exp(-\frac{u_{i}^{i}+u_{i}^{c}+\sum_{j}^{j\in\mathcal{N}_{i}}(1.5u_{j}^{c}-0.5u_{j}^{n})-C}{4})))$ (13)
where $C$ is a constant set to ensure the positivity assumption holds. And then, $z$ can be obtained by aggregating $t$ using $E$.
Sequentially, the outcome $y$ is generated by
$y_{i}=u^{i}_{i}+u^{c}_{i}+t_{i}\times(2u^{i}+1.6u^{c})+z_{i}\times(\sum_{j}^{j\in\mathcal{N}_{i}}(1.5u_{j}^{c}-0.5u_{j}^{n})+1.5u^{i}_{i}+0.5u^{c}_{i})+\epsilon_{y,i},$ (14)
where $\epsilon_{y,i}$ is the Gaussian noise.
Finally, we generate the proxies $x$ of latent confounders by $x_{i}=w[u^{i}_{i},u^{c}_{i},u^{n}_{i}]+\epsilon_{x,i}$ where $w$ is a randomly $3\times 6$ weight matrix sampled from Gaussian distribution with a mean $1$ and scale $1$, and $\epsilon_{x,i}$ is the Gaussian noise. This results in the final $6$-dimensional observed proxies $x$.
C.3 Additional Experimental Results
We report additional experimental results on BC(homo) and BC(hete) datasets in Table 4 and Table 5. The results are similar to the experimental results in the Flickr dataset.
Table 4: Experimental results on BC(homo) Dataset. The top result is highlighted in bold, and the runner-up is underlined.
$\varepsilon_{average}$ $\varepsilon_{individual}$
Within Sample Out-of Sample Within Sample Out-of Sample
Methods AME ASE ATE AME ASE ATE IME ISE ITE IME ISE ITE
CFR+z $0.0788_{\pm 0.0096}$ $0.1157_{\pm 0.0076}$ $0.2323_{\pm 0.0106}$ $0.0770_{\pm 0.0099}$ $0.1157_{\pm 0.0075}$ $0.2306_{\pm 0.0106}$ $0.0796_{\pm 0.0091}$ $0.1158_{\pm 0.0076}$ $0.2325_{\pm 0.0106}$ $0.1000_{\pm 0.0388}$ $0.1157_{\pm 0.0075}$ $0.2405_{\pm 0.0166}$
ND+z $0.2375_{\pm 0.0450}$ $0.0316_{\pm 0.0104}$ $0.0790_{\pm 0.0226}$ $0.2380_{\pm 0.0458}$ $0.0323_{\pm 0.0122}$ $0.0768_{\pm 0.0254}$ $0.2377_{\pm 0.0448}$ $0.0321_{\pm 0.0101}$ $0.0792_{\pm 0.0226}$ $0.2477_{\pm 0.0460}$ $0.0379_{\pm 0.0099}$ $0.1068_{\pm 0.0172}$
NetEst $0.1059_{\pm 0.0609}$ $0.0284_{\pm 0.0297}$ $0.0387_{\pm 0.0288}$ $0.0987_{\pm 0.0663}$ $0.0257_{\pm 0.0276}$ $0.0356_{\pm 0.0268}$ $0.1366_{\pm 0.0481}$ $0.0631_{\pm 0.0205}$ $0.0994_{\pm 0.0214}$ $0.1680_{\pm 0.0620}$ $0.0920_{\pm 0.0316}$ $0.1507_{\pm 0.0647}$
TNet $0.1045_{\pm 0.0610}$ $0.0502_{\pm 0.0559}$ $0.0473_{\pm 0.0229}$ $0.1045_{\pm 0.0610}$ $0.0502_{\pm 0.0559}$ $0.0473_{\pm 0.0229}$ $0.1045_{\pm 0.0610}$ $0.0502_{\pm 0.0559}$ $0.0473_{\pm 0.0229}$ $0.1045_{\pm 0.0610}$ $0.0502_{\pm 0.0559}$ $0.0473_{\pm 0.0229}$
Ours_w/o_IPM $0.0356_{\pm 0.0176}$ $0.0222_{\pm 0.0104}$ $0.0514_{\pm 0.0163}$ $0.0416_{\pm 0.0176}$ $0.0206_{\pm 0.0100}$ $0.0433_{\pm 0.0136}$ $0.0411_{\pm 0.0147}$ $0.0244_{\pm 0.0122}$ $0.0554_{\pm 0.0182}$ $0.0435_{\pm 0.0169}$ $0.0218_{\pm 0.0106}$ $0.0453_{\pm 0.0137}$
Ours $0.0661_{\pm 0.0485}$ $0.0232_{\pm 0.0164}$ $0.0442_{\pm 0.0258}$ $0.0661_{\pm 0.0485}$ $0.0232_{\pm 0.0164}$ $0.0442_{\pm 0.0258}$ $0.0661_{\pm 0.0485}$ $0.0232_{\pm 0.0164}$ $0.0442_{\pm 0.0258}$ $0.0661_{\pm 0.0485}$ $0.0232_{\pm 0.0164}$ $0.0442_{\pm 0.0258}$
Table 5: Experimental results on BC(hete) Dataset. The top result is highlighted in bold, and the runner-up is underlined.
$\varepsilon_{average}$ $\varepsilon_{individual}$
Within Sample Out-of Sample Within Sample Out-of Sample
Methods AME ASE ATE AME ASE ATE IME ISE ITE IME ISE ITE
ND+z $0.4124_{\pm 0.0702}$ $0.0451_{\pm 0.0201}$ $0.1330_{\pm 0.0205}$ $0.4111_{\pm 0.0737}$ $0.0486_{\pm 0.0206}$ $0.1326_{\pm 0.0261}$ $0.4226_{\pm 0.0673}$ $0.0562_{\pm 0.0146}$ $0.1941_{\pm 0.0246}$ $0.4330_{\pm 0.0662}$ $0.0666_{\pm 0.0137}$ $0.2211_{\pm 0.0321}$
NetEst $0.1643_{\pm 0.1337}$ $0.0450_{\pm 0.0180}$ $0.0594_{\pm 0.0262}$ $0.1857_{\pm 0.1168}$ $0.0405_{\pm 0.0252}$ $0.0343_{\pm 0.0179}$ $0.2199_{\pm 0.1039}$ $0.0667_{\pm 0.0171}$ $0.1731_{\pm 0.0368}$ $1.5595_{\pm 2.5329}$ $1.1347_{\pm 1.9028}$ $1.0924_{\pm 1.7278}$
TNet $0.1216_{\pm 0.0864}$ $0.0537_{\pm 0.0524}$ $0.0429_{\pm 0.0301}$ $0.1257_{\pm 0.0727}$ $0.0537_{\pm 0.0511}$ $0.0481_{\pm 0.0269}$ $0.1731_{\pm 0.0450}$ $0.0655_{\pm 0.0465}$ $0.1458_{\pm 0.0175}$ $0.1915_{\pm 0.0542}$ $0.0650_{\pm 0.0458}$ $0.1740_{\pm 0.0621}$
Ours_w/o_IPM $0.0706_{\pm 0.0324}$ $0.0278_{\pm 0.0079}$ $0.0605_{\pm 0.0504}$ $0.0755_{\pm 0.0420}$ $0.0294_{\pm 0.0052}$ $0.0664_{\pm 0.0457}$ $0.1240_{\pm 0.0237}$ $0.0491_{\pm 0.0060}$ $0.1706_{\pm 0.0311}$ $0.1273_{\pm 0.0276}$ $0.0481_{\pm 0.0054}$ $0.1705_{\pm 0.0269}$
Ours $0.0625_{\pm 0.0597}$ $0.0195_{\pm 0.0162}$ $0.0579_{\pm 0.0255}$ $0.0604_{\pm 0.0610}$ $0.0195_{\pm 0.0130}$ $0.0598_{\pm 0.0259}$ $0.1175_{\pm 0.0380}$ $0.0373_{\pm 0.0099}$ $0.1545_{\pm 0.0141}$ $0.1172_{\pm 0.0392}$ $0.0365_{\pm 0.0074}$ $0.1563_{\pm 0.0130}$

### v2

@@ VI-B1 RQ1: Effectiveness of CaLaNet
As shown in Tables I, II, III, and IV, we have conducted experiments by running our proposed method, CaLaNet, and several baselines. Overall, our CaLaNet outperforms all methods consistently with smaller estimation errors in all metrics, indicating the effectiveness of our methods. Specifically, compared with the baselines, our methods perform better in terms of both average and heterogeneous treatment effect estimation, with smaller errors and standard deviation. This means our CaLaNet is not only accurate for networked effect estimation but also performs stably. This is reasonable since most existing methods do not consider the latent confounders that hinder the effect identification, while our method utilizes representation learning techniques and thereby achieves superior performances with recovered latent confounders. In particular, SPNet+z also considers latent confounders. However, it has no theoretical guarantee for the latent confounder recovery, and thus underperforms our CaLaNet. Hence, the results in III, Table I, II, and IV showcase the effectiveness of our proposed CaLaNet.
Figure 4: The figure presents scatter plots visualizing the relationship between the recovered latent confounders and ground-truth latent confounders $U^{i},U^{c}$, and $U^{n}$.
@@
TABLE V: MCC results between the recovered latent confounders and ground-truth latent confounders $U^{i},U^{c}$, and $U^{n}$.
$\hat{U}^{i}$ $\hat{U}^{c}$ $\hat{U}^{n}$
${U}^{i}$ 0.9729 0.1742 0.2032
${U}^{c}$ 0.3796 0.9090 0.3616
${U}^{n}$ 0.1879 0.2420 0.9901
VI-B2 RQ2: Correctness of Representation Learning
To validate the correctness of our representation learning method, we conduct experiments in the simulated dataset and visualize the recovered latent confounders $\hat{U}^{i},\hat{U}^{c},\hat{U}^{n}$ alongside ground-truth ones $U^{i},U^{c},U^{n}$ in Figure 4. We also report the corresponding Matthews Correlation Coefficient (MCC) results in Table V. As shown in Figure 4, the diagonal scatter plots (top-left, middle, and bottom-right) show strong alignments, indicating that each recovered latent confounder ($\hat{U}^{i},\hat{U}^{c},\hat{U}^{n}$) closely corresponds to its ground-truth counterpart ($U^{i},U^{c},U^{n}$). In contrast, the off-diagonal scatter plots exhibit no obvious correlation, suggesting that the latent factors are well disentangled. Similar findings are reflected in the MCC results in Table V, where the corresponding recovered and ground-truth confounders show high correlation, while the non-corresponding ones do not. These results confirm that the proposed representation learning method effectively recovers the latent confounders, in line with the theoretical identifiability guarantees.
Figure 5: Hyperparameter sensitivity result regarding individual effect estimation error on BC(homo) dataset.
Figure 6: Hyperparameter sensitivity result regarding individual effect estimation error on Flickr(homo) dataset.
VI-B3 RQ3: Stability Regarding Hyperparameters
[…]
VII Conclusion
In this paper, we tackle the challenge of identifying and estimating causal effects under networked interference in the presence of latent confounders. Specifically, we systematically categorize three distinct types of latent confounders that hinder networked effect identification. To address this issue, we propose leveraging the networked information to achieve the identifiability of latent confounders. Based on this recovery, we theoretically establish conditions under which networked causal effects become identifiable. Building on these theoretical insights, we develop an effective estimator grounded in identifiable representation learning. Extensive experiments empirically validate both the theoretical results and the practical effectiveness of our proposed method.
Acknowledgments
This research was supported in part by National Science and Technology Major Project (2021ZD0111501), National Science Fund for Excellent Young Scholars (62122022), Natural Science Foundation of China (U24A20233, 62206064, 62206061, 62476163, 62406078), Guangdong Basic and Applied Basic Research Foundation (2023B1515120020), and CCF-DiDi GAIA Collaborative Research Funds (CCF-DiDi GAIA 202311).
VIII References Section